An apartment has the dimensions 20 m by 10 m by 7 m. The temperature is 25°C, and the relative humidity is 51 percent. What is the total mass (in kg) of water vapor in the air in the apartment?


1
Expert's answer
2021-08-02T08:39:21-0400

Given:

"V=20\times 10\times 7=1400\:\rm m^3"

"\varphi=51\%"

"\rho_{\rm sv}=23\:\rm g/m^3"

The relative humidity

"\varphi=\frac{\rho}{\rho_{\rm sv}}\times 100\%"

The density of water vapor in the air

"\rho=\rho_{\rm sv}*\frac{\varphi}{100\%}=23*\frac{51\%}{100\%}=11.73\:\rm g/m^3"

The total mass (in kg) of water vapor in the air in the apartment

"m=\rho*V=11.73*1400=16422\:\rm g=16.422\:\rm kg"